# Generating Verifiable Chain of Thoughts from Exection-Traces

### v2

@@ 5.2. Impact of Data Curation
For the Granite-3.3-8b model, the results are definitive. As shown in Table 2, the 25k correctness-filtered dataset substantially outperforms the others across all benchmarks. On LiveCodeBench, it achieves a score of 44.9%, an absolute gain of +26.6 over the base model, and significantly outperforms the larger 54k set (34.1%). This powerful improvement from a targeted, high-quality dataset indicates that for complex reasoning, data verifiability is far more impactful than sheer volume.
When repeating the experiment with Qwen2.5-Coder-7B, all three fine-tuned models dramatically outperform the base model. For instance, the 25k dataset boosts the LiveCodeBench score from 46.3% to 67.0% (+20.7). Unlike with Granite, the performance across the three data subsets is competitive. The 25k dataset achieves the highest score on LiveCodeBench, while the 54k dataset has a slight edge on the CruxEval benchmarks. Given the consistently strong performance of the 25k set, we selected it as the winning configuration for all subsequent experiments.
Figure 7. Performance comparison against SOTA baselines (ding2024semcodertrainingcodelanguage) Our model’s results are highlighted with a hatch pattern. Abbreviations: CL: CodeLlama, SC2: StarCoder2, DSCoder: DeepSeekCoder, MCoder: MagicCoder, -Py: Python, -Inst: Instruct.
5.3. Impact of Reasoning Direction
Using the winning 25k dataset from the previous stage, we evaluated the impact of our novel bi-directional data format. We compared models trained on forward-only, backward-only, and the complete bi-directional datasets.
For the Granite-3.3-8b model, the results in Table 2 highlight the benefits of bi-directional training. The model trained on the bi-directional dataset achieves the highest scores on both CruxEval Output (45.7% pass@1) and Input (42.1% pass@1). This represents a final performance gain of +30.2 on CruxEval Output over the base model, demonstrating that teaching cause-to-effect and effect-to-cause reasoning provides a synergistic improvement.
This finding is further confirmed with the Qwen2.5-Coder-7B model. The bi-directionally trained model once again emerges as the top performer, achieving the best results on LiveCodeBench (68.2%) and CruxEval Input (61.9% pass@1). For this configuration, final score on LiveCodeBench represents a +21.9 point gain over the base model. To contextualize this performance, Figure 7 compares our best model against a range of competitive foundation models. Our fine-tuned model establishes a leading score on LiveCodeBench-Execution and demonstrates highly competitive performance across all reasoning benchmarks.
5.4. Does the method used to generate and verify chain-of-thought rationales fundamentally impact model reasoning capabilities, or are all CoT approaches equivalent?
To answer this question, we systematically evaluate how the source and validation of rationales affect downstream model performance. We are not only comparing verification strategies, but the different approaches to CoT generation: execution-grounded rationales, LLM-generated rationales (potentially hallucinated), and no rationales at all.
Using the same 25k problem set from our data curation ablation, we generated training data under six distinct conditions and fine-tuned Qwen2.5-Coder-7B models for comparison:
  • Baseline Approaches: (1) Base Model: No fine-tuning; (2) Few-Shot: 3-shot prompting with example problems, no fine-tuning.
  • Alternative CoT Strategies: (3) Q&A Only: Fine-tuned on question-answer pairs with no CoT rationales; (4) LLM-Generated CoT: Fine-tuned on CoT rationales generated purely by an LLM without execution grounding.
  • Our Execution-Grounded Approach: (5) Minimal Execution: Our trace-grounded CoT using minimal verification (1 solution, 1 test); (6) Full Approach: Our complete pipeline with Dual Agreement verification (10 solutions, 25 tests) and trace-grounded rationales.
Table 3 presents the complete results. The ablation reveals a clear hierarchy of components, with three key insights. First, execution grounding is the dominant factor: LLM-generated CoT, despite providing step-by-step rationales, fails because it trains models on plausible but unverified reasoning patterns. This validates our core thesis that the source of rationales—whether grounded in actual execution or generated from statistical patterns—determines whether models internalize verifiable semantics or learn superficial imitation. Second, verification quality compounds this effect: even when using execution-grounded rationales, insufficient candidate diversity degrades performance, as single solution-test pairs cannot filter edge-case failures. This confirms that consensus through Dual Agreement is not merely a scalability mechanism but a necessary correctness guarantee. Third, the step-by-step reasoning itself is essential: removing CoT rationales entirely causes the largest performance drop, demonstrating that models require explicit traces of how outputs arise from inputs, not just input-output correspondences. Prompting alone provides negligible improvement, underscoring that these reasoning capabilities must be internalized through fine-tuning on verified data rather than retrieved at inference time.
@@ 5.3. Impact of Reasoning Direction
Table 3. Impact of verification approach on fine-tuning results. All models fine-tuned on 25k bi-directional data with different verification methods. Tested on Qwen2.5-Coder-7B.
Verification Approach LiveCodeBench CruxEval-O CruxEval-I HumanEval
Base Model (No FT) 46.3 45.3 47.5 62.0
Few-Shot (3 examples) 48.1 47.0 48.8 63.5
Q&A Only (No CoT) 52.7 50.2 51.3 69.0
LLM-as-Judge + LLM CoT 59.3 54.1 55.7 75.5
Our Approach (10S/25T + Trace) 68.2 59.7 61.9 81.5
5.5. Enhancing Instruction-Tuned Models
@@
To confirm that our best-performing bi-directional data provides a specialized signal complementary to general instruction tuning, we fine-tuned existing instruct-tuned models. As shown in Figure 8, this provides a substantial boost, yielding a sharp +39.9 gain on CruxEval Input for granite-3.3-8B-instruct and a +21.5 gain on CruxEval Output for the already strong Qwen2.5-Coder-7B-Instruct, demonstrating the value of our execution-grounded data.
@@
Figure 8. Performance boost from fine-tuning instruct models. Solid bars represent the baseline instruct models; hatched bars show the improvement after fine-tuning with the best-performing 25k bi-directional subset of data.
5.6. Chain-of-Thought Consistency and Quality Analysis
To assess the quality of reasoning generated by our fine-tuned model, we compare CoTs generated by our fine-tuned model Qwen2.5-Coder-7B-FT against the base Qwen2.5-Coder-7B model on CruxEval-O problems. For each problem, we generate CoT from both models and evaluate two key dimensions: consistency between reasoning and final answer, and information content of the reasoning content.
Consistency Analysis. We measure consistency using a composite metric that evaluates reasoning-answer alignment across multiple dimensions: entailment patterns, conceptual overlap, and sequential coherence. Figure 9 (top) shows our model exhibits strong correlation between CoT length and consistency ($R^{2}=0.122$), indicating longer reasoning is more logically coherent. The baseline model shows no such relationship ($R^{2}=0.011$). Moreover, when CoT is consistent with the final answer, our model is more likely to produce correct solutions (AUC = 0.567) compared to the baseline where consistency poorly predicts correctness (AUC = 0.502).
Information Richness. The bottom plot reveals that our model produces substantially more informative reasoning. Using an entropy-based metric that accounts for vocabulary diversity and technical term density, we find the CoTs from our fine-tuned model are 761% more information-rich than those from the baseline. This difference is statistically significant with a very large effect size (Cohen’s $d=7.93$, $p<0.001$), confirming that our training method generates more detailed and semantically meaningful rationales.
These results confirm that verified CoT training fundamentally enhances reasoning quality, producing more consistent and informative chain-of-thought processes that better support final outputs.
Figure 9. Chain-of-Thought quality analysis. Top: CoT-to-outcome consistency vs reasoning length with regression analysis. Bottom: Information content distributions. Verified CoT training shows superior consistency scaling ($R^{2}=0.122$ vs $0.011$) and 761% higher information richness ($d=7.93$, $p<0.001$).
6. Conclusion and Future Work
We introduced a methodology for generating verifiable CoT data for code reasoning, addressing the challenge of logical integrity in synthetic datasets. By grounding every reasoning step in program execution traces, our pipeline eliminates logical hallucinations and produces high-fidelity training data that is correct by construction.
Our systematic evaluation confirms the effectiveness of this approach. We show that verified data quality is more important than quantity and that our novel bi-directional format significantly improves reasoning. These findings hold across different model architectures, where fine-tuning boosted performance by as much as +30.2 points on the CruxEval-O reasoning benchmark, underscoring the fundamental impact of our contribution.
Future work includes extending our language-agnostic pipeline to other languages like C++ and Java and exploring its use in advanced training schemes such as offline reinforcement learning with DPO. To aid the development of more robust language models, we will publicly release our complete synthesis pipeline.
@@ 6. Conclusion and Future Work
This appendix provides concrete examples of the verifiable Chain-of-Thought (CoT) data generated by our synthesis pipeline. The samples are organized to showcase the diversity of our dataset across two key dimensions: reasoning direction and problem difficulty. We include examples for:
  • Reasoning Direction: Forward-only, Backward-only, and Bi-directional conversational formats.
  • Problem Difficulty: Samples rated at levels 2 (Easy), 3 (Medium), and 4 (Hard) based on our quality-scoring mechanism.
Each sample is presented in a structured format that distinguishes the user prompt (the problem context) from the model’s generated response (the verifiable reasoning and final answer).
@@
Table 4. SFT Hyper-parameters

### v3

@@ 5.1 Impact of Data Curation
For Granite-3.3-8B, the 25k correctness-filtered dataset substantially outperforms others. On LCB, it achieves 44.9%, an absolute gain of +26.6 over the base model, well ahead of the larger 54k set (34.1%). This indicates that for complex reasoning, data verifiability is far more impactful than volume.
For Qwen2.5-Coder-7B, all fine-tuned models outperform the baseline. While the 54k set edges ahead on CruxEval, the 25k dataset achieves the best LCB score (67.0% Pass@1, +20.7) with comparable CruxEval performance. Given its stronger execution reasoning and smaller training cost, we selected the 25k set for subsequent experiments. Figure 6 compares our results against published baselines.
5.2 Impact of Reasoning Direction
Using the 25k dataset, we evaluated forward-only, backward-only, and bi-directional training. Table 2 shows that bi-directional training improves both CruxEval and LiveCodeBench performance.
For Granite-3.3-8B, the bi-directional dataset achieves the highest scores on CruxEval Output (45.7% Pass@1, +30.2) and Input (42.1% Pass@1, +27.8), while maintaining near-best LCB performance (44.3 vs. 44.9 for forward-only). This demonstrates that teaching both cause-to-effect and effect-to-cause reasoning yields complementary gains on understanding tasks without sacrificing execution performance.
For Qwen2.5-Coder-7B, bi-directional training again achieves the best results on LiveCodeBench (68.2%) and CruxEval Input (61.9%), representing +21.9 on LCB over the base model.
5.3 Impact of Verification Approach
We ablate verification methods using the 25k problem set with six conditions: (1) Base Model, (2) Few-Shot prompting, (3) Q&A Only without CoT, (4) LLM-Generated CoT verified by an LLM judge (Mistral-Small-24B) without execution grounding, (5) Minimal Execution (1 solution, 1 test), and (6) Our Full Approach with Dual Agreement (Table 3). The LLM-Judge + LLM CoT condition is our closest non-execution-grounded multi-candidate baseline. For each problem, we generate multiple CoT candidates and use a judge model to select the best rationale-and-answer pair. While this differs from majority-vote self-consistency, it captures the same intuition of improving quality through candidate selection without access to execution traces.
@@ 5.3 Impact of Verification Approach
Table 3: Impact of verification approach (Pass@1). The table isolates the contribution of CoT supervision, execution grounding, and consensus-based artifact selection. All fine-tuned models use 25k bi-directional data. Tested on Qwen2.5-Coder-7B.
Approach LCB CruxEval O/I HumanEval
Base Model 46.3 45.3/47.5 62.0
Few-Shot (3-shot) 48.1 47.0/48.8 63.5
Q&A Only (no CoT) 52.7 50.2/51.3 69.0
LLM-Judge + LLM CoT 59.3 54.1/55.7 75.5
Min. Exec (1s, 1t) 62.1 55.8/57.2 77.0
Ours (Full) 68.2 59.7/61.9 81.5
[…]
5.4 Enhancing Instruction-Tuned Models
@@
To confirm our execution-grounded data provides complementary benefits to instruction tuning, we also fine-tuned instruct-tuned versions of the models. As shown in Figure 7, this yields a +39.9 gain on CruxEval Input for Granite-3.3-8B-Instruct and a +21.5 gain on CruxEval Output for Qwen2.5-Coder-7B-Instruct.
@@
Figure 7: Performance boost from fine-tuning instruct models. Solid bars: baseline; hatched bars: improvement after fine-tuning with 25k bi-directional data.
5.5 CoT Quality Analysis
In addition to task accuracy, we ask whether execution-grounded training also improves the quality of the reasoning itself. We compare CoTs generated by fine-tuned Qwen2.5-Coder-7B against the baseline on CruxEval Output along two dimensions.
Consistency. We measure how well reasoning steps align with the final answer. Our model shows a much stronger relationship between reasoning detail and logical correctness ($R^{2}=0.122$ vs. baseline $0.011$), and when its reasoning is internally consistent, it produces correct answers more reliably (AUC =0.567=0.567 vs. $0.502$). In other words, our model’s reasoning chains are not just longer—they are more logically coherent.
Information richness. We measure how much technical detail each explanation contains using Shannon entropy over the distribution of unique technical tokens (variable names, concrete values, and control flow keywords) per explanation—higher entropy indicates more diverse, specific content. Compared to the base model, our fine-tuned model produces CoTs that are 761% richer in information content ($p<0.001$). It references specific variable values, traces state transitions, and describes actual control flow from the execution—rather than generating generic summaries.
5.6 Deployment Footprint
The pipeline has been released as open source and integrated into downstream model post-training workflows. In particular, our 25k verified CoT subset was used in post-training stages of Granite-family of models IBM Research (2025), and the pipeline has also been adapted for use with standard open fine-tuning stacks such as OpenInstruct OLMo et al. (2025). These integrations demonstrate practical downstream use as a reusable systems contribution that has already been exercised in real training pipelines.
6 Conclusion
We presented a method to generate factually grounded CoT training data at scale by grounding each reasoning step in execution traces. Results demonstrate that verified data quality drives improved reasoning, yielding peak gains of up to +22.2 on CruxEval, +26.6 on LiveCodeBench-Exec, and +19.5 on HumanEval over the corresponding base models we fine-tune. Beyond open-sourcing the pipeline, the resulting verified CoT data has already been used in post-training stages of deployed Granite-family models and integrated with existing fine-tuning stacks, indicating practical downstream utility. Future work includes extending the trace-acquisition layer to additional languages such as Java while preserving the same verification framework.
@@ 6 Conclusion
  • Evaluation Setup — Benchmark descriptions and official links
  • Experimental Setup — Compute infrastructure and SFT hyperparameters
  • Dual Agreement Consensus Validation — Empirical validation and hyperparameter selection
  • Data Generation Prompts — Prompts for instruction, signature, code, test, and CoT narration
  • Dual Agreement Clustering Sample — Concrete pipeline example
  • Verifiable CoT Data Samples — Examples across reasoning directions and difficulty levels
  • Trace-to-CoT Mapping — End-to-end example from execution trace to bi-directional CoT
  • Sliding Window Verification — Algorithm details and rejection example
@@
Table 4: SFT Hyper-parameters
Dual Agreement Consensus Validation
Figure 8: Percentage of problems achieving high consensus $\text{Score}(C_{i})$ across configurations of candidate solutions ($m$, x-axis) and candidate tests ($n$, y-axis). Both dimensions contribute: $m$ strengthens cluster size $|C_{i}|$ while $n$ improves test coverage $|T_{p}(C_{i})|$.
We validate the Dual Agreement verification approach through a controlled study on 5,000 programming problems with ground-truth solutions. For each configuration of $m$ candidate solutions and $n$ candidate tests, we compute clusters $C_{i}$ and their quality scores $\text{Score}(C_{i})=|C_{i}|\times|T_{p}(C_{i})|$ as defined in Section 3.2, then measure the percentage of problems where the highest-scoring cluster achieves a high consensus score.
Figure 8 shows these consensus rates as a function of $m$ and $n$. The bottom-right region (high $m$, high $n$) achieves 70–90% of problems with high consensus scores, while the upper-left (low $m$, low $n$) shows near-zero rates. This validates that both dimensions are necessary: increasing $m$ provides more independent solutions for stronger cluster agreement ($|C_{i}|$), while increasing $n$ provides broader test coverage ($|T_{p}(C_{i})|$), both of which contribute to higher $\text{Score}(C_{i})$.
Score Formula Validation
Figure 9: Left: Cluster size $|C_{i}|$ vs. tests passed $|T_{p}(C_{i})|$ for the top-scoring cluster of each problem, colored by correctness. Iso-score curves show decision boundaries at $\text{Score}=15,30,60$. Right: Correctness rate by score range with 95% Wilson confidence intervals. Higher $\text{Score}(C_{i})=|C_{i}|\times|T_{p}(C_{i})|$ yields significantly higher correctness ($p<0.01$), confirming the formula reliably selects correct solutions.
Figure 10: Hyperparameter selection for Dual Agreement. Left: Alignment with ground truth across configurations of candidate solutions ($m$) and tests ($n$). Quality plateaus beyond 25 tests. Right: Quality-cost tradeoff showing the $5\times 25$ configuration achieves $\sim$90% of maximum quality at 12.5% of the computational cost.
Figure 8 shows that consensus improves with more candidates, but does the scoring formula $\text{Score}(C_{i})=|C_{i}|\times|T_{p}(C_{i})|$ actually distinguish correct solutions from incorrect ones? Figure 9 answers this directly.
The left panel plots each problem’s top cluster in the $|C_{i}|$ vs. $|T_{p}(C_{i})|$ space, colored by whether the selected solution matches the ground truth. Correct solutions (green) concentrate in the upper-right region where both cluster size and test coverage are high, while incorrect solutions (red) scatter in the lower-left. The iso-score curves ($\text{Score}=15,30,60$) show that the multiplicative structure of the formula naturally creates decision boundaries—problems above the $\text{Score}=30$ curve are almost always correct.
The right panel quantifies this relationship by binning problems into score ranges and computing correctness rates with 95% Wilson confidence intervals. The results show a clear monotonic relationship: problems with $\text{Score}<15$ have only 6.1% correctness, while those with $\text{Score}>60$ achieve 96.9% correctness ($p<0.01$ between low and mid score ranges). The error bars confirm that these differences are statistically significant, not artifacts of sampling variance. This validates the formula’s design: a solution must have both strong agreement among independent implementations ($|C_{i}|$) and broad test coverage ($|T_{p}(C_{i})|$) to be selected, making it statistically unlikely that an incorrect solution survives.
Candidate Generation Hyperparameters: Quality-Cost Analysis
While Figure 8 above measures consensus—whether independently generated solutions agree with each other—it does not directly tell us whether the selected solution is actually correct. High consensus means many solutions behave identically, but they could all be wrong in the same way.
To address this, we conduct a separate analysis measuring alignment with ground truth: for each configuration of $m$ candidate solutions and $n$ candidate tests, we run Dual Agreement on a validation set of 500 problems with known-correct solutions, and measure how often the selected cluster’s canonical solution matches the ground-truth answer.
Figure 10 presents this analysis in two panels:
Left panel shows alignment with ground truth as a heatmap across configurations. Unlike the consensus heatmap, which measures internal agreement, this measures external correctness. The key observations are: (1) quality generally increases with both the number of candidate solutions and tests, though with natural variance across individual cells; (2) a clear plateau emerges at approximately 25 tests—beyond this point, adding more tests yields marginal improvements regardless of the number of solutions; and (3) the $5\times 25$ configuration (marked with a star) achieves 84.9% alignment, which falls at the entry of the plateau region.
Right panel plots the same data as a quality-cost tradeoff curve, where cost is measured as the total number of solution-test executions ($m\times n$). This view makes the diminishing returns explicit: the $5\times 25$ configuration requires only 125 executions and achieves approximately 90% of the maximum quality obtained by the most expensive $20\times 50$ configuration (1,000 executions). The shaded region highlights the zone of diminishing returns where substantially higher computational cost yields only marginal quality gains.
In summary, the two heatmaps are complementary: Figure 8 validates that Dual Agreement produces reliable consensus (an internal property), while Figure 10 confirms that this consensus translates to actual correctness (an external property) and identifies the optimal operating point at $5\times 25$.